# On the Reachability of 3–Dimensional Paths with a Prescribed Curvature Bound

### v2

@@ 3.2.1 Implicit description of the reachability set
Now, we will show that the endpoint $(\boldsymbol{x}(t_{f}),\boldsymbol{y}(t_{f}))$ lies interior to $\mathcal{G}_{3}(t_{f})$ if $N>3$ in case D. Suppose the corresponding $\boldsymbol{x}(t)$ of $\boldsymbol{\Xi}$ in case D is planar, consists of multiple C’s with more than three components, denoted by $C_{1}C_{2}C_{3}\dots C_{N-1}C_{N}$. According to the conclusion in case D, $C_{2}$ and $C_{3}$ have same lengths $<2\pi$, and $C_{1}$ has a length not greater than the two. Thus, the sum of the lengths of $C_{1}$ and $C_{3}$ is larger than $C_{2}$ and none of them contains a cycle. Consequently, Lemma 2 in Patsko et al. (2003) asserts that there exists an alternative CCCCC curve, planar, such that its length, initial and terminal locations and directions coincide with $C_{1}C_{2}C_{3}$. Furthermore, it was proved that the lengths of the third and fourth C’s are different in such CCCCC curves. By replacing the $C_{1}C_{2}C_{3}$ component by such curve, the resulting $CCCCCC_{4}\dots C_{N-1}C_{N}$ curve has identical initial and terminal locations and directions with $C_{1}C_{2}C_{3}\dots C_{N-1}C_{N}$. If this alternative trajectory is another outcome of case D, then the middle C’s must have same lengths. However, this is not true. But along the conclusions of every other cases among A to E, we find that such trajectory is not a possible outcome of the PMP conditions. Therefore, the endpoints of the trajectories in case D with $N>3$ have alternative trajectories reaching the point that only allow trivial adjoint response. Consequently, they cannot lie on the boundary of $\mathcal{G}_{3}(t_{f})$.
Up to this point are the proof of items (1) and (2) of this theorem. Next, we prove the remaining items by identifying certain trajectories that posses an alternative trajectory reaching the same state endpoint $(\boldsymbol{x}(t_{f}),\boldsymbol{y}(t_{f}))$, but only allows trivial adjoint response.
@@
Figure 3: Illustration of auxiliary trajectories that only allow trivial adjoint response. Among the state variables, only the $\boldsymbol{x}$ component is depicted.
@@
Lastly, we observe from part 2 that the CCC trajectories necessary for constructing the boundary of $\mathcal{G}_{3}(t_{f})$ arise only from case D. This is because the CCC trajectories from case E can be reduced to CC and hence are redundant. Then, the conclusion of case D asserts that all C components have lengths $<2\pi$. This completes the proof of the remaining items (3), (4), and (5). ∎ Items (1) and (2) of Theorem 11 collectively serve as the counterpart to Theorem 1 in Patsko et al. (2003) under 3D setting. Items (3), (4), and (5) of Theorem 11 are 3D counterparts of the developments in Buzikov and Galyaev (2022) (Lemma 4 and Theorem 6). It should be noted that Theorem 1 in Patsko et al. (2003) as well as Theorem 11 in this paper, additional conditions to eliminate redundant trajectories with endpoints in the interior of reachability sets are required for complete determination of the boundary. For 2D curves, Patsko and Fedotov (2022) provided such additional conditions and enabled complete determination of $\mathcal{G}_{2}(t_{f})$. We leave the development of the counterpart of Patsko and Fedotov (2022) for 3D curves as future work.
@@ 3.2.2 Position reachability set
In the remaining part of this section, we present the complete description of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$, or the position reachability set. The results expand the 2D case studied in Cockayne and Hall (1975). In this case, unlike Theorem 11, complete determination is possible by dint of additional geometric statements made in Cockayne and Hall (1975). Let us commence by stating the below theorem, which follows from the proof of Theorem 11.
@@
In case D, a lifted controlled arc $\boldsymbol{\Xi}$ satisfying the PMP conditions consists of concatenation of finite number of nice trajectories, denoted as $\boldsymbol{\Xi}_{1}\boldsymbol{\Xi}_{2}\dots\boldsymbol{\Xi}_{N}$. Moreover, $\boldsymbol{\Xi}_{i}$ for $12\leq i\leq N-$ 1 correspond to a maximal solution of Eq. (22) and therefore have the same lengths. But $\boldsymbol{W}(t_{f})=0$ further indicates that $\boldsymbol{\Xi}_{N}$ also corresponds to a maximal solution. Therefore, all $\boldsymbol{\Xi}_{i}$’s, except possibly $\boldsymbol{\Xi}_{1}$, must have the same lengths. Additionally, since the endpoints $(\boldsymbol{x}(t_{f}),\boldsymbol{y}(t_{f}))$ of trajectories in $N>3$ cases lie interior to $\mathcal{G}_{3}(t_{f})$, their projections by $\pi_{3}$ also lie interior to $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$, given that projection is an open map. In other words, $N$ is at most 3 for $\boldsymbol{x}(t_{f})$ to lie on the boundary of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$. If $N=3$ and if we write $C_{1}C_{2}C_{3}$, the lengths of $C_{2}$ and $C_{3}$ are the same. Consequently, Lemma 2 in Patsko et al. (2003) implies that there exists an alternative CCCCC trajectory. During the proof of Theorem 11, we have seen that this cannot be an outcome of the PMP conditions. Hence, trajectories rising from case D required for construction of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ are confined to C or planar CC trajectories such that the length of the second C is not less than the first. Additionally, the CC trajectories naturally have length $\leq 4\pi$ because each nice trajectory is defined on intervals of length $<2\pi$.
In Case E of the proof of Theorem 11, it is proved that each of the nice trajectories, corresponding to the C components, is of length $2\pi$ and rises from a maximal solution of Eq. (22), provided that its domain does not include the initial time 00 or the final time $t_{f}$. However, if $\boldsymbol{W}(t_{f})=0$, the last C component corresponds to a maximal nice trajectory as well, and its length is $2\pi$. Consequently, the same process as in the proof of case E in Theorem 11 can be applied: such cycle can be reattached tangentially to the initial point. This reduces the overall curve into curves of CS or its subsegments. This completes the proof. ∎ It should be noted that as in part 3 of the proof of Theorem 11, Theorem 12 can be strengthened through identifying alternative trajectories that only allow trivial adjoint response. However, we do not bother such tuning because the geometric developments in Cockayne and Hall (1975) are sufficient for complete determination of $\pi_{3}\left(\mathcal{G}(t_{f})\right)$.
@@
To facilitate further discussion, let us introduce additional notations. From here on, we will denote by $\ell$ the line in $\mathbb{R}^{3}$ passing through the point $\boldsymbol{x}(0)$ with direction $\boldsymbol{y}(0)$. An arbitrary plane in $\mathbb{R}^{3}$ containing $\ell$ is indicated as $P$. Recall that $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ represents the set of all endpoints of curvature bounded paths in $\mathbb{R}^{2}$. Henceforth, under the same curvature bound setting, we abuse the notation and indicate by $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ the set of all endpoints of planar curves that lie entirely in $P$. In other words, the ambient space is considered as $P$ instead of $\mathbb{R}^{2}$. Schematic illustration of boundaries of $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ in $\mathbb{R}^{2}$ and in $P$ are depicted in Fig. 4.
@@
To apply these conclusions to determine $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$, let us denote by $E_{2}^{P}(t_{f})$ the set of endpoints in $P$ of all planar curves in the classes CS, CC, and their subsegments. The difference with $E_{1}^{P}(t_{f})$ is that the curves need not be entirely contained in $P$ but are only required to be planar. Since $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ is invariant under rotation around the axis $\ell$, it can be generated by rotating $\mathcal{S}(t_{f})\equiv\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)\cap P$ around $\ell$. It is easy to see that $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)\subseteq\mathcal{S}(t_{f})$ because the responses of Eq. (9) include the planar curves that lie in $P$. Moreover, $\mathcal{S}(t_{f})$ varies continuously in time if and only if $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ does. In the followings, we will show that $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)=\mathcal{S}(t_{f})$ by using continuity of $\mathcal{S}(t_{f})$.
It follows from Theorem 12 that $E_{2}^{P}(t_{f})$ contains the boundary of $\mathcal{S}(t_{f})$, and the union of $E_{2}^{P}(t_{f})$ over all $P$ contains the boundary of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$. We introduce the following lemma.
@@
Proof. From definition, it is evident that $E_{1}^{P}(t_{f})\subseteq E_{2}^{P}(t_{f})$. We will show the inclusion relationship in the converse direction. Suppose a point $\boldsymbol{z}$ in $E_{2}^{P}(t_{f})$ is reached by a planar curve in classes CS, CC, or their subsegments that lies entirely in a plane $P^{\prime}$. It follows from definition that $\boldsymbol{z}$ is in both $P$ and $P^{\prime}$. Since $\ell$ is the tangent line of the curve at $t=0$, it follows $\ell\subseteq P^{\prime}$. The claim is trivial if $P=P^{\prime}$, so we assume $P\neq P^{\prime}$. Then $\ell$ is a line entirely contained in two distinct planes $P$ and $P^{\prime}$. Thus, $\ell=P\cap P^{\prime}$ and $\boldsymbol{z}\in\ell$. Then one can rotate the entire curve—which originally lies in $P^{\prime}$—around $\ell$ until the curve lies in $P$. This maintains the endpoint on $\boldsymbol{z}$ because it is on the axis of rotation, $\ell$. The result ing curve is of length $t_{f}$, lies in $P$, belongs to class CS, CC, or their subsegments, and reaches $\boldsymbol{z}$. This proves that $\boldsymbol{z}\in E_{1}^{P}(t_{f})$. ∎ Now, the points in $E_{1}^{P}(t_{f})$—or equivalently, $E_{2}^{P}(t_{f})$—that violate any of the five items in Proposition 7 lie interior to $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ and, consequently, interior to $\mathcal{S}(t_{f})$ since $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)\subseteq\mathcal{S}(t_{f})$. Thus, Theorem 12 can be strengthened as follows: $E_{2}^{P}(t_{f})\setminus E^{\prime P}(t_{f})$ contains the boundary of $\mathcal{S}(t_{f})$. But given $E_{1}^{P}(t_{f})\setminus E^{\prime P}(t_{f})$, or equivalently $E_{2}^{P}(t_{f})\setminus E^{\prime P}(t_{f})$, there is a unique way to choose the enclosed regions in $P$ such that their union varies continuously in time. This union equals to $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ and $\mathcal{S}(t_{f})$. This proves that $\mathcal{S}(t_{f})=\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ and, subsequently, the following theorem.
@@
Theorem 14 allows complete determination of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ for all $t_{f}$, as depicted in Fig. 5.
@@
Figure 5: Cross–sectional view of the evolution of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ over multiple values of $t_{f}$.
@@
(Remark on the general dimensional case)
The structure of extremal trajectories for the Markov–Dubins problem in general dimensions was studied in Mittenhuber (1998), also based on PMP. The analysis extended beyond euclidean spaces to include noneuclidean spaces such as spheres, $\mathbb{S}^{n}$, and hyperbolic spaces, $\mathbb{H}^{n}$. It was proved that to solve the Markov–Dubins problem in $\mathbb{R}^{n}$ (resp. $\mathbb{S}^{n}$, $\mathbb{H}^{n}$), it suffices to solve it in $\mathbb{R}^{3}$ (resp. $\mathbb{S}^{3}$, $\mathbb{H}^{3}$). This suggests a promising avenue for applying the approach of this paper to derive more general conclusions on reachability in general dimensions. However, this requires to entirely reformulate the problem—including the conclusions on reachability in Theorems 3 and 4—on the Lie group $SE(n)$ of Euclidean motions of $\mathbb{R}^{n}$. We leave this extension for future work.

### v3

@@ 3.2.1 Implicit description of the reachability set
Now, we will show that the endpoint $(\boldsymbol{x}(t_{f}),\boldsymbol{y}(t_{f}))$ lies interior to $\mathcal{G}_{3}(t_{f})$ if $N>3$ in case D. Suppose the corresponding $\boldsymbol{x}(t)$ of $\boldsymbol{\Xi}$ in case D is planar, consists of multiple C’s with more than three components, denoted by $C_{1}C_{2}C_{3}\dots C_{N-1}C_{N}$. According to the conclusion in case D, $C_{2}$ and $C_{3}$ have same lengths $<2\pi$, and $C_{1}$ has a length not greater than the two. Thus, the sum of the lengths of $C_{1}$ and $C_{3}$ is larger than $C_{2}$ and none of them contains a cycle. Consequently, Lemma 2 in Patsko et al. (2003) asserts that there exists an alternative CCCCC curve, planar, such that its length, initial and terminal locations and directions coincide with $C_{1}C_{2}C_{3}$. Furthermore, it was proved that the lengths of the third and fourth C’s are different in such CCCCC curves. By replacing the $C_{1}C_{2}C_{3}$ component by such curve, the resulting $CCCCCC_{4}\dots C_{N-1}C_{N}$ curve has identical initial and terminal locations and directions with $C_{1}C_{2}C_{3}\dots C_{N-1}C_{N}$. If this alternative trajectory is another outcome of case D, then the middle C’s must have same lengths. However, this is not true. But along the conclusions of every other cases among A to E, we find that such trajectory is not a possible outcome of the PMP conditions. Therefore, the endpoints of the trajectories in case D with $N>3$ have alternative trajectories reaching the same endpoint by a controller that is not maximal. Consequently, they cannot lie on the boundary of $\mathcal{G}_{3}(t_{f})$. This completes the proof of items (1) and (2). Next, we prove the remaining items by identifying certain trajectories that posses an alternative trajectory reaching the same state endpoint $(\boldsymbol{x}(t_{f}),\boldsymbol{y}(t_{f}))$ by a controller that is not maximal.
@@
Figure 3: Illustration of auxiliary trajectories that are generated from controllers that are not maximal. Among the state variables, only the $\boldsymbol{x}$ component is depicted.
@@
Lastly, we observe from part 2 that the CCC trajectories necessary for constructing the boundary of $\mathcal{G}_{3}(t_{f})$ arise only from case D. This is because the CCC trajectories from case E can be reduced to CC and hence are redundant. Then, the conclusion of case D asserts that all C components have lengths $<2\pi$. This completes the proof of the remaining items (3), (4), and (5). ∎ Items (1) and (2) of Theorem 8 collectively serve as the counterpart to Theorem 1 in Patsko et al. (2003) in 3D. Items (3), (4), and (5) of Theorem 8 are 3D counterparts of the developments in Buzikov and Galyaev (2022) (Lemma 4 and Theorem 6). It should be noted that Theorem 1 in Patsko et al. (2003) and Theorem 8 in this paper, additional conditions to eliminate redundant trajectories with endpoints in the interior of reachability sets are required for complete determination of the boundary. For 2D curves, Patsko and Fedotov (2022) provided such additional conditions and enabled complete determination of $\mathcal{G}_{2}(t_{f})$. We leave the development of the counterpart of Patsko and Fedotov (2022) for 3D curves as future work.
@@ 3.2.2 Position reachability set
Finally, we present the complete description of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$, or the position reachability set. The results expand the 2D case studied in Cockayne and Hall (1975). Unlike Theorem 8, complete determination is done by additional geometric statements. We commence by stating the below theorem.
@@
In case D, a lifted controlled arc $\boldsymbol{\Xi}$ satisfying the PMP conditions consists of concatenation of finite number of nice trajectories, denoted as $\boldsymbol{\Xi}_{1}\boldsymbol{\Xi}_{2}\dots\boldsymbol{\Xi}_{N}$. Moreover, $\boldsymbol{\Xi}_{i}$ for $12\leq i\leq N-$ 1 correspond to a maximal solution of Eq. (22) and therefore have the same lengths. But $\boldsymbol{W}(t_{f})=0$ further indicates that $\boldsymbol{\Xi}_{N}$ also corresponds to a maximal solution. Therefore, all $\boldsymbol{\Xi}_{i}$’s, except possibly $\boldsymbol{\Xi}_{1}$, must have the same lengths. Additionally, since the endpoints $(\boldsymbol{x}(t_{f}),\boldsymbol{y}(t_{f}))$ of trajectories in $N>3$ cases lie interior to $\mathcal{G}_{3}(t_{f})$, their projections by $\pi_{3}$ also lie interior to $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$, given that projection is an open map. In other words, $N$ is at most 3 for $\boldsymbol{x}(t_{f})$ to lie on the boundary of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$. If $N=3$ and if we write $C_{1}C_{2}C_{3}$, the lengths of $C_{2}$ and $C_{3}$ are the same. Consequently, Lemma 2 in Patsko et al. (2003) implies that there exists an alternative CCCCC trajectory. During the proof of Theorem 8, we confirmed that this cannot be an outcome of the PMP conditions. Hence, trajectories rising from case D required for construction of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ are confined to C or planar CC trajectories such that the length of the second C is not less than the first.
@@
In Case E of the proof of Theorem 8, it is proved that each of the nice trajectories, corresponding to the C components, is of length $2\pi$ and rises from a maximal solution of Eq. (22), provided that its domain does not include the initial time 00 or the final time $t_{f}$. However, if $\boldsymbol{W}(t_{f})=0$, the last C component corresponds to a maximal nice trajectory as well, and its length is $2\pi$. Consequently, the same process as in the proof of case E in Theorem 8 can be applied: such cycle can be reattached tangentially to the initial point. This reduces the overall curve into curves of CS or its subsegments. ∎ From here on, we denote by $\ell$ the line in $\mathbb{R}^{3}$ passing through the point $\boldsymbol{x}(0)$ with direction $\boldsymbol{y}(0)$. An arbitrary plane in $\mathbb{R}^{3}$ containing $\ell$ is indicated as $P$. Recall that $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ represents the set of all endpoints of curvature bounded paths in $\mathbb{R}^{2}$. Henceforth, under the same curvature bound setting, we abuse the notation and indicate by $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ the set of all endpoints of planar curves that lie entirely in $P$. In other words, the ambient space is considered as $P$ instead of $\mathbb{R}^{2}$. Schematic illustration of boundaries of $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ in $\mathbb{R}^{2}$ and in $P$ are depicted in Fig. 4.
@@
Now, let us denote by $E_{2}^{P}(t_{f})$ the set of endpoints in $P$ of all planar curves in the classes CS, CC, and their subsegments. The difference with $E_{1}^{P}(t_{f})$ is that the curves need not be entirely contained in $P$ but are only required to be planar. Since $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ is invariant under rotation around the axis $\ell$, it can be generated by rotating $\mathcal{S}(t_{f})\equiv\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)\cap P$ around $\ell$. It is easy to see that $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)\subseteq\mathcal{S}(t_{f})$ because the responses of Eq. (9) include the planar curves that lie in $P$. Moreover, $\mathcal{S}(t_{f})$ varies continuously in time if and only if $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ does.
Subsequently, we show that $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)=\mathcal{S}(t_{f})$. It follows from Theorem 9 that $E_{2}^{P}(t_{f})$ contains the boundary of $\mathcal{S}(t_{f})$, and the union of $E_{2}^{P}(t_{f})$ over all $P$ contains the boundary of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$. We introduce the following lemma.
@@
Proof. From definition, it is evident that $E_{1}^{P}(t_{f})\subseteq E_{2}^{P}(t_{f})$. We will show the inclusion relationship in the converse direction. Suppose a point $\boldsymbol{z}$ in $E_{2}^{P}(t_{f})$ is reached by a planar curve in classes CS, CC, or their subsegments that lies entirely in a plane $P^{\prime}$. It follows from definition that $\boldsymbol{z}$ is in both $P$ and $P^{\prime}$. Since $\ell$ is the tangent line of the curve at $t=0$, it follows $\ell\subseteq P^{\prime}$. The claim is trivial if $P=P^{\prime}$, so we assume $P\neq P^{\prime}$. Then $\ell$ is a line entirely contained in two distinct planes $P$ and $P^{\prime}$. Thus, $\ell=P\cap P^{\prime}$ and $\boldsymbol{z}\in\ell$. Then one can rotate the entire curve—which originally lies in $P^{\prime}$—around $\ell$ until the curve lies in $P$. This maintains the endpoint on $\boldsymbol{z}$ because it is on the axis of rotation, $\ell$. The result ing curve is of length $t_{f}$, lies in $P$, belongs to class CS, CC, or their subsegments, and reaches $\boldsymbol{z}$. This proves that $\boldsymbol{z}\in E_{1}^{P}(t_{f})$. ∎ Now, the points in $E_{1}^{P}(t_{f})$—or equivalently, $E_{2}^{P}(t_{f})$—that violate any of the five items in Proposition 7 lie interior to $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ and, consequently, interior to $\mathcal{S}(t_{f})$ since $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)\subseteq\mathcal{S}(t_{f})$. This strengthens Theorem 9 as follows: $E_{2}^{P}(t_{f})\setminus E^{\prime P}(t_{f})$ contains the boundary of $\mathcal{S}(t_{f})$. But given $E_{1}^{P}(t_{f})\setminus E^{\prime P}(t_{f})$, or equivalently $E_{2}^{P}(t_{f})\setminus E^{\prime P}(t_{f})$, there is a unique way to choose the enclosed regions in $P$ such that their union varies continuously in time. This union equals to $\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ and $\mathcal{S}(t_{f})$. This proves that $\mathcal{S}(t_{f})=\pi_{2}\left(\mathcal{G}_{2}(t_{f})\right)$ and, subsequently, the following theorem. This allows complete determination of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ for all $t_{f}$ as in Fig. 5.
@@
Figure 5: Cross-sectional view of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$ over multiple $t_{f}$.
@@
(Remark on the general-dimensional case) The structure of maximal trajectories for the Markov-Dubins problem in general dimensions was studied in Mittenhuber (1998), also based on PMP. The analysis extended beyond euclidean spaces to include noneuclidean spaces such as spheres, $\mathbb{S}^{n}$, and hyperbolic spaces, $\mathbb{H}^{n}$. It was proved that to solve the Markov-Dubins problem in $\mathbb{R}^{n}$ (resp. $\mathbb{S}^{n}$, $\mathbb{H}^{n}$), it suffices to solve it in $\mathbb{R}^{3}$ (resp. $\mathbb{S}^{3}$, $\mathbb{H}^{3}$). This suggests a promising avenue for applying the approach of this paper to derive conclusions in general dimensions. However, this requires to entirely reformulate the problem on the Lie group $SE(n)$ of Euclidean motions of $\mathbb{R}^{n}$. We leave this extension for future work.